OpenStudy (anonymous):

Draw a line that has the given slope and y- intercept. 7x=5y

OpenStudy (anonymous):

slope = coefficient of x y intercept is where x equals zero

OpenStudy (anonymous):

Hhhmm still confused as to how I would write it.

OpenStudy (anonymous):

Should it look something like this? y=7x+0

OpenStudy (anonymous):

well in has to be in linear form so isolate y first 7/5x=y or y = 7/5x therefore your slope would be 7/5

OpenStudy (anonymous):

ok so you are dividing 5 by both sides to get rid of y right?

OpenStudy (anonymous):

So now I understand how you got 7/5x but now as far as the line how would that appear. I am so lost and this really confuses me. Sorry

OpenStudy (anonymous):

it's alright =) write out a x and y intercepts chart and you can see point such as x=5 y=7 x=10 y=14 x=15 y=21 and then you just graph them and draw a line through all the points

OpenStudy (anonymous):

May I ask another question?

OpenStudy (anonymous):

Write an equation of the line containing the specified point and perpendicular to the indicated line. (-3, 7) Y=5

OpenStudy (anonymous):

I came up with -1/5x+ 4/5 is this correct or did I do it correctly

OpenStudy (anonymous):

Hello? Are you still there?

OpenStudy (anonymous):

yea hold on; it's 3/5 for the last variable I believe other than that everything looks great!

OpenStudy (anonymous):

Ok instead of 4/5 it should be 3/5?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

I am not sure what it is that I did wrong. y-y1=m(x-x1)

OpenStudy (anonymous):

y-(-2)= -1/5(x-4)

OpenStudy (anonymous):

yea that's the correct form; i suppose you just substituted 4 instead of -3 before multiplying it by -1/5

OpenStudy (anonymous):

the coordinates are (-3, 7)...

OpenStudy (anonymous):

I came up with y+2= -1/5(x-4)

OpenStudy (anonymous):

when I multiply -1/5 x -4 i got the answer 0.8

OpenStudy (anonymous):

which in fraction I got 4/5 ....is this correct

OpenStudy (anonymous):

I am all confused now bec now when I subtract -2 from each side I get -1/2

OpenStudy (anonymous):

never mind that was the incorrect equation let me try the correct one which is (4,-2) x=1

OpenStudy (anonymous):

How would I solve for this?